OpenStudy (anonymous):

2(x + 6) − (x + 1) = 16

OpenStudy (anonymous):

2x+12-x-1 =16 2x-x =16-12+1 proceed

OpenStudy (anonymous):

im not sure how to get the numbers to the other side

OpenStudy (anonymous):

but you can open up the parenthesis?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so we can proceed from here: 2x+12-x-1 =16

OpenStudy (anonymous):

you add like terms so x+11=16 ?

OpenStudy (anonymous):

right

OpenStudy (anonymous):

cause they're negatives

OpenStudy (anonymous):

okay thank you

OpenStudy (anonymous):

you welcome